OpenStudy (anonymous):

What is 20% of 25% of 4/5? a) 0.0025 b) 0.004 c) 0.005 d) 0.04 e) 0.05

OpenStudy (mathstudent55):

0.2 * 0.25 * 0.8 = ?

OpenStudy (anonymous):

1/5*1/4*4/5=1/25

OpenStudy (anonymous):

0.04i believe the aanswer would be. :)

OpenStudy (anonymous):

Thank you! @mathstudent55 and @hahd !